# VegasAfterglow: A High-Performance Framework for Gamma-Ray Burst Afterglows

### v1

@@ 3.2.2 Reverse Shock Crossing via energy conservation
Figure 5 shows the shock crossing radius coefficient $C_{\Delta}$ as a function of magnetization $\sigma$ for different values of $\Gamma_{43}-1$. Consistent with the result in Zhang and Kobayashi (2005), we find that for $\Gamma_{43}>1.5$, $C_{\Delta}$ can be approximated as $\frac{3}{4}(1+\sigma)^{-1/2}$. However, in the thin-shell regime, where the reverse shock is weak and $\Gamma_{43}-1\ll 1$ at early times, $C_{\Delta}$ begins to deviate from this scaling well before $\sigma\sim 1$. As indicated in the figure, $C_{\Delta}$ starts to decrease significantly at lower values of $\sigma$, leading to the discrepancy between our numerical results and the analytical scaling in Figure 4. The late-time flattening observed in Figure 5 arises from the use of the exact expression for the relative Lorentz factor (Eq. 33), rather than the commonly used approximation $\Gamma_{43}\sim\frac{1}{2}\left(\frac{\Gamma_{4}}{\Gamma_{3}}+\frac{\Gamma_{3}}{\Gamma_{4}}\right)$, which may break down if either $\Gamma_{4}$ or $\Gamma_{3}$ does not satisfy the $\gg 1$ condition.
@@ 3.2.3 Post Shock Crossing Dynamics
Once the reverse shock crosses the entire jet, the reverse shock will gradually adjust to the Blandford-McKee self-similar solution (Blandford and McKee, 1976), while maintaining the shocked proton number constant afterward. The forward shock will continue propagating through the external medium and the Lorentz factor of the forward shock will be determined by the blast wave equation (Chiang and Dermer, 1999; Piran, 1999; Huang et al., 1999; Pe’er, 2012; Nava et al., 2013; Zhang, 2018).
@@
Here, we carefully take the pressure, radiation and adiabatic expansion of the blast wave into consideration. From the stress energy tensor, one can write the blast wave equation as (Zhang, 2018),
${d}[\Gamma_{2}M_{\rm tot}c^{2}+\Gamma_{\rm eff}U]$ $=$ $c^{2}{dm}+\Gamma_{\rm eff}dU_{\rm rad}+L_{\rm inj}dt_{\rm eng}\,,$ (56)
${dU}$ $=$ ${dU_{\rm sh}}+dU_{\rm ad}+{dU_{\rm rad}}\,,$ (57)
$M_{\rm tot}$ $=$ $M_{0}+m+m_{\rm inj}\,,$ (58)
where $E_{\rm jet}=\Gamma_{2}M_{0}c^{2}$ with $M_{0}$ being the equivalent initial rest mass of the jet, $m$ is the total collected mass defined in Section 2.3, $m_{\rm inj}$ is the injected mass, $U$ is the internal energy in the blast wave, $dU_{\rm sh}$ is increased internal energy due to shock heating, $dU_{\rm ad}$ is the lost internal energy due to adiabatic expansion, $dU_{\rm rad}$ is lost internal energy due to radiation, and $L_{\rm inj}$ is the injection luminosity. Those quantities are defined as
$\Gamma_{\rm eff}$ $=$ $\frac{\hat{\gamma}_{2}\Gamma_{2}^{2}-\hat{\gamma}_{2}+1}{\Gamma_{2}}\,,$ (59)
$dU_{\rm sh}$ $=$ $c^{2}(\Gamma_{2}-1)dm\,,$ (60)
$dU_{\rm ad}$ $=$ $-(\hat{\gamma}_{2}-1)\bigg{(}\frac{3}{r}dr-\frac{1}{\Gamma_{2}}d\Gamma_{2}\bigg{)}U\,,$ (61)
$dU_{\rm rad}$ $=$ $-\epsilon(\Gamma_{2}-1)c^{2}dm\,,$ (62)
$\epsilon$ $=$ $\epsilon_{\rm rad}\epsilon_{e}\,.$ (63)
The radiation efficiency writes
$\epsilon_{\rm rad}=\left\{\begin{aligned} &1,\quad&\gamma_{c}<\gamma_{m}&\\ &\bigg{(}\frac{\gamma_{m}}{\gamma_{c}}\bigg{)}^{2-p},\quad&\gamma_{c}>\gamma_{m}&\end{aligned}\right.\,,$ (64)
for fast cooling and slow cooling, respectively (See definition of $\gamma_{m}$ and $\gamma_{c}$ in Section 4).
The blast wave equation then for $\Gamma_{2}$ and $U$ can be written as
$\dot{\Gamma}_{2}$ $=$ $\frac{-(\Gamma_{\rm eff}+1)(\Gamma_{2}-1)c^{2}\dot{m}-\Gamma_{\rm eff}\dot{U}_{\rm ad}+L_{\rm inj}-c^{2}\Gamma_{2}\dot{m}_{\rm inj}}{(M_{0}+m+m_{\rm inj})c^{2}+U\frac{d\Gamma_{\rm eff}}{d\Gamma_{2}}}\,,$ (65)
$\dot{U}$ $=$ $(1-\epsilon)(\Gamma_{2}-1)c^{2}\dot{m}-(\hat{\gamma}_{2}-1)\bigg{(}\frac{3}{r}\dot{r}-\frac{\dot{\Gamma}_{2}}{\Gamma_{2}}\bigg{)}U\,,$ (66)
where [^3] One may note that $\dot{U}_{\rm ad}$ is also a function of $\dot{\Gamma}_{2}$, so one may need to move the corresponding terms in $\dot{U}_{\rm ad}$ to the left side of the equation to get the final expression of $\dot{\Gamma}_{2}$.
$\dot{r}=\frac{dr}{dt}=\frac{c\beta_{2}}{1-\beta_{2}}\,.$ (67)
Then $n_{2}$ and $B_{2}$ can be obtained by using Equation 14 and 31. One may note that setting $\hat{\gamma}_{2}=1$ (i.e., ignoring the pressure in the blast wave) does not recover the equation for $\dot{\Gamma}_{2}$ from Huang et al. (1999). This discrepancy arises because of different assumptions about shock heating. In Huang et al. (1999), the authors assume $dU_{\rm sh}=c^{2}d[(\Gamma_{2}-1)m]$, whereas here we adopt $dU_{\rm sh}=c^{2}(\Gamma_{2}-1)dm$.
(i) Reverse shock
For the reverse shock, the scaling laws shortly after the reverse shock crossing cannot be solved analytically. Mészáros and Rees (1999) introduced a parameterized power-law decay behavior where
@@
This treatment automatically recovers all the regimes with correct $\Delta_{3}^{\prime}$ spreading where,
$\Delta_{3}^{\prime}$ $=$ $\Delta^{\prime}_{\times}+{\rm c_{s}}t^{\prime}\,,$ (76)
$c_{s}^{2}$ $=$ $c^{2}{\frac{\hat{\gamma}_{43}p}{\rho_{0}c^{2}+\frac{\hat{\gamma}_{43}}{\hat{\gamma}_{43}-1}p}}=c^{2}{\frac{\hat{\gamma}_{43}(\hat{\gamma}_{43}-1)(\Gamma_{43}-1)}{1+\hat{\gamma}_{43}(\Gamma_{43}-1)}}\,,$ (77)
with $t^{\prime}$ being the co-moving time. The shell starts to expand adiabatically after the shock crossing radius, $r_{\times}$,
$\Delta_{3}^{\prime}\propto{\rm c_{s}}\frac{r}{\Gamma_{3}}\propto\left\{\begin{aligned} &\sqrt{\hat{\gamma}_{43}\frac{p_{3}}{n_{3}}}\frac{r}{\Gamma_{3}}\propto r^{\frac{2(2+g-\hat{\gamma}_{43})}{\hat{\gamma}_{43}+1}},\quad&\Gamma_{43}-1\ll 1\\ &\sqrt{\hat{\gamma}_{43}-1}\frac{r}{\Gamma_{3}}\propto r^{1+g},\quad&\Gamma_{43}-1\gg 1\end{aligned}\right.\,.$ (78)
@@ 4.1 Minimal injection Lorentz factor
In the context of relativistic shocks, the dissipation of kinetic energy at the shock front leads to a sharp increase in internal energy and pressure in the downstream region, accelerating particles to relativistic energies. The comoving internal energy density behind the shock can be obtained from Equation 66
@@
where $\gamma_{m}$ and $\gamma_{M}$ (See definition in Section 4.3) are the minimum and maximum Lorentz factor of electrons, respectively. The minimum Lorentz factor, $\gamma_{m}$, can then be determined from Eqs. 90 and 91, see, e.g. Sari et al. 1998; Panaitescu and Kumar 2002; Dai and Cheng 2001; Zhang 2018, but with $(+1)$ added at the end:
@@
[…]
The $+1$ correction on $\gamma_{m}$ comes from that the kinetic electron energy $\epsilon_{e}\mathcal{U}\propto\gamma_{m}-1$ (not $\propto\gamma_{m}$). This is valid in both relativistic and non-relativistic regimes and guarantees that $\gamma_{m}>1$. However, at low electron energies, synchrotron radiation transitions into the cyclotron regime, where emission is dominated by the fundamental harmonic, and relativistic broadening becomes negligible. Since the electron population follows a power-law distribution in energy, a portion of the electrons may lie close to or within the cyclotron regime, where synchrotron emission is inefficient. To account for this, a correction factor, $f_{\rm syn}$, is required to suppress the synchrotron flux at lower energies, defined as
@@
$\frac{dN_{e}}{dx}\propto(x^{2}+2x)^{-(p+1)/2}(x+1)\propto\gamma(\gamma^{2}+1)^{-(p+1)/2}\,,$ (95)
where $x=\gamma-1$. With this distribution, one can find
@@
We remark that, without the $+1$ correction on $\gamma_{m}$ in the non-relativistic regime, the flux in the deep Newtonian regime will be underestimated. Furthermore, with the $+1$ correction on $\gamma_{m}$ but without the $f_{\rm syn}$ correction, the flux in the deep Newtonian regime will be overestimated (overflattened).
@@ 4.2 Synchrotron & IC Cooling
$\frac{d\gamma}{dt^{\prime}}m_{e}c^{2}=-(P_{\rm syn}+P_{\rm IC})=-\frac{4}{3}\sigma_{T}c\gamma^{2}\beta^{2}\frac{B^{\prime 2}}{8\pi}(1+\tilde{Y}),$ (97)
@@ 4.4 Synchrotron Self-absorption
where $I_{\nu}$ is the specific synchrotron intensity (see section 4.6) and $kT$ is the effective blackbody temperature. Then one can solve the self-absorption frequency for weak absorption $\nu_{a}<{\rm min}(\nu_{m},\nu_{c})$ (intersection in the $1/3$ segment), intermediate absorption $\nu_{m}<\nu_{a}<\nu_{c}$ (intersection in the $-(p-1)/2$ segment) and strong absorption $\nu_{c}<\nu_{a}$ (electron pile-up around $\gamma_{a}$)
@@
where $\Sigma_{\rm p,tot}=\frac{N_{\rm p,tot}}{4\pi r^{2}}$ is the column density of the shocked proton and $N_{\rm p,tot}$ is the total number of shocked protons. The coefficient $0.92$ arises from the maximum value of the single-electron synchrotron spectrum function $\mathcal{F}(x)$ (Rybicki and Lightman, 1979), and the factor $\frac{\pi}{4}$ comes from the average pitch angle $\langle\sin\alpha\rangle$.
@@ 4.6 Synchrotron spectrum
Fig. 8 shows an example of on-axis observed synchrotron spectra for narrow opening angle top-hat jets (thus mimicking the shape of the intrinsic spectrum) with afterglow parameters in the six different regimes.
@@ 4.7 Inverse Compton Cooling
where below $\gamma_{c}$ ($\nu_{c}$), the electron (synchrotron) spectrum remains unchanged, while above $\gamma_{c}$ ($\nu_{c}$), a correction factor $\propto 1/(1+\tilde{Y})$ has been applied.
@@
for slow cooling. The broken power-law/broken power-law division in Eqs. 119 and 120 give rise to additional spectral segments above $\gamma_{c}$ ($\nu_{c}$). As shown in Figure 8, the blue lines correspond to synchrotron-only cooling, the green lines to Thomson IC cooling with constant $\tilde{Y}$, and the orange lines to spectra including Klein–Nishina corrections, which introduce additional breaks for $\nu>\nu_{c}$.
@@ 4.8 Inverse Compton Spectrum
[…]
@@ 5 Geometry
[…]
@@ 6 Examples
Medium type ISM ISM/Wind ISM/Wind ISM/Wind Arbitrary
@@ 6.1 Strctured Jets
First, we present examples of structured jets observed from various viewing angles and compare them with the results from AfterglowPy and Jetsimpy. The left panel of Figure 10 shows the light curve for a top-hat jet with a half-opening angle of $5^{\circ}$. All three codes agree closely in this simple case; however, AfterglowPy exhibits a slight deviation at small viewing angles during the early coasting phase due to its approximate treatment of that regime. The middle panel illustrates a Gaussian jet with a $5^{\circ}$ core angle, while the right panel shows a power-law jet with index 2 and the same core angle. For structured jets, both AfterglowPy and Jetsimpy predict higher fluxes. This arises from their different treatments of the angular Lorentz‐factor profile: AfterglowPy allows $\Gamma(\theta)$ to fall below unity but truncates emission beyond a “wing” angle $\theta_{\rm wing}$, whereas Jetsimpy and VegasAfterglow model $\Gamma-1$ with a Gaussian profile and apply the “+1” correction.

### v3

@@ 3.2.2 Reverse Shock Crossing via energy conservation
Figure 5 shows the shock crossing radius coefficient $C_{\Delta}$ as a function of magnetization $\sigma$ for different values of $\Gamma_{43}-1$. Consistent with the result in Zhang and Kobayashi (2005), we find that for $\Gamma_{43}>1.5$, $C_{\Delta}$ can be approximated as $\frac{3}{4}(1+\sigma)^{-1/2}$. However, in the thin-shell regime, where the reverse shock is weak and $\Gamma_{43}-1\ll 1$ at early times, $C_{\Delta}$ begins to deviate from this scaling well before $\sigma\sim 1$. As indicated in the figure, $C_{\Delta}$ starts to decrease significantly at lower values of $\sigma$, leading to the discrepancy between our numerical results and the analytical scaling in Figure 4. The late-time flattening observed in Figure 5 arises from the use of the exact expression for the relative Lorentz factor (Equation 33), rather than the commonly used approximation $\Gamma_{43}\sim\frac{1}{2}\left(\frac{\Gamma_{4}}{\Gamma_{3}}+\frac{\Gamma_{3}}{\Gamma_{4}}\right)$, which may break down if either $\Gamma_{4}$ or $\Gamma_{3}$ does not satisfy the $\gg 1$ condition.
We remark that the reverse shock generation condition given in Equation 43 remains valid in the energy conservation approach, since it is derived at the onset of the reverse shock crossing, when the blast wave is still infinitesimally thin. Under this circumstance, pressure balance and energy conservation are indeed equivalent.
@@ 3.2.3 Post Shock Crossing Dynamics
Once the reverse shock crosses the entire jet, it will gradually adjust to the Blandford–McKee self-similar solution (Blandford and McKee, 1976), while maintaining a constant number of shocked protons afterward. The forward shock will continue to propagate through the external medium. Several early works have addressed the blast wave dynamics after the reverse shock crossing (Chiang and Dermer, 1999; Piran, 1999; Huang et al., 1999; Pe’er, 2012; Nava et al., 2013; Zhang, 2018; Yuan et al., 2024).
@@
After the shock crossing, the shock heating and mass accumulation in region 3 cease, which we implement by turning off the corresponding terms in Equation 61, where
$\frac{dm_{3}}{dt}$ $=$ $0,$ (66)
$\frac{dU_{3}}{dt}$ $=$ $-(\hat{\gamma}_{34}-1)U_{3}\bigg(\frac{2}{r}\frac{dr}{dt}+\frac{1}{\Delta_{3}^{\prime}}\frac{d\Delta_{3}^{\prime}}{dt}\bigg),$ (67)
$\frac{d\Delta_{3}^{\prime}}{dt}$ $=$ $c_{\rm s,3}\frac{dt^{\prime}}{dt},$ (68)
$c_{s,3}$ $=$ $c\sqrt{\frac{\hat{\gamma}_{34}(\hat{\gamma}_{34}-1)(\Gamma_{3}-1)}{1+\hat{\gamma}_{34}(\Gamma_{3}-1)}}.$ (69)
One can then continue solving Equation 61 to obtain the forward shock dynamics. Indeed, if the adiabatic cooling in region 3 is ignored (cold blast wave assumption), this reduces to the forward shock dynamics given in Equation 8.60 of Zhang (2018). It is worth noting that further simplification by setting $\hat{\gamma}_{2}=1$ (i.e., neglecting the pressure in the blast wave) does not recover the shock dynamics in Huang et al. (1999). This discrepancy arises from different assumptions about shock heating. In Huang et al. (1999), the authors adopt $dU_{\rm sh}=c^{2}d[(\Gamma_{2}-1)m_{2}]$, whereas here we adopt $dU_{\rm sh}=c^{2}(\Gamma_{2}-1)dm_{2}$.
(ii) Reverse shock
After the reverse shock has crossed the shell, the assumption $\Gamma_{2}=\Gamma_{3}$ is no longer valid, as regions 2 and 3 subsequently evolve with generally different bulk Lorentz factors. The scaling laws shortly after the reverse shock crossing cannot be solved analytically. Mészáros and Rees (1999) introduced a parameterized power-law decay behavior where
@@
This treatment automatically recovers all the regimes with correct $\Delta_{3}^{\prime}$ spreading. The shell starts to expand adiabatically after the shock crossing radius, $r_{\times}$,
$\Delta_{3}^{\prime}\propto{\rm c_{s,3}}\frac{r}{\Gamma_{3}}\propto\left\{\begin{aligned} &\sqrt{\hat{\gamma}_{43}\frac{p_{3}}{n_{3}}}\frac{r}{\Gamma_{3}}\propto r^{\frac{2(2+g-\hat{\gamma}_{43})}{\hat{\gamma}_{43}+1}},\quad&\Gamma_{43}-1\ll 1\\ &\sqrt{\hat{\gamma}_{43}-1}\frac{r}{\Gamma_{3}}\propto r^{1+g},\quad&\Gamma_{43}-1\gg 1\end{aligned}\right.\,$. (78)
@@ 4.1 Minimal injection Lorentz factor
In the context of relativistic shocks, the dissipation of kinetic energy at the shock front leads to a sharp increase in internal energy and pressure in the downstream region, accelerating particles to relativistic energies. The comoving internal energy density behind the shock can be obtained from
@@
where $\gamma_{m}$ and $\gamma_{M}$ (See definition in Section 4.3) are the minimum and maximum Lorentz factor of electrons, respectively. The minimum Lorentz factor, $\gamma_{m}$, can then be determined from Equation 90 and 91, see, e.g. Sari et al. 1998; Panaitescu and Kumar 2002; Dai and Cheng 2001; Zhang 2018, but with $(+1)$ added at the end:
@@
$\bar{\gamma}_{m}-1=\frac{\int_{\gamma_{m}}^{\gamma_{M}}(\gamma-1)\frac{dN_{e}}{d\gamma}d\gamma}{\int_{\gamma_{m}}^{\gamma_{M}}\frac{dN_{e}}{d\gamma}d\gamma}\,.$ (93)
The +1+1 correction on $\gamma_{m}$ comes from that the kinetic electron energy $\epsilon_{e}\mathcal{U}\propto\gamma_{m}-1$ (not $\propto\gamma_{m}$). This is valid in both relativistic and non-relativistic regimes and guarantees that $\gamma_{m}>1$. However, at low electron energies, synchrotron radiation transitions into the cyclotron regime, where emission is dominated by the fundamental harmonic, and relativistic broadening becomes negligible. Since the electron population follows a power-law distribution in energy, a portion of the electrons may lie close to or within the cyclotron regime, where synchrotron emission is inefficient. To account for this, a correction factor, $f_{\rm syn}$, is required to suppress the synchrotron flux at lower energies, defined as
@@
$\frac{dN_{e}}{dy}\propto(y^{2}+2y)^{-(p+1)/2}(y+1)\propto\gamma(\gamma^{2}+1)^{-(p+1)/2}\,$, (95)
where $y=\gamma-1$. With this distribution, one can find
@@
We remark that, without the +1+1 correction on $\gamma_{m}$ in the non-relativistic regime, the flux in the deep Newtonian regime will be underestimated. Furthermore, with the +1+1 correction on $\gamma_{m}$ but without the $f_{\rm syn}$ correction, the flux in the deep Newtonian regime will be overestimated (overflattened).
@@ 4.2 Synchrotron & IC Cooling
[…]
@@ 4.4 Synchrotron Self-absorption
where $I_{\nu}$ is the specific synchrotron intensity (see section 4.6) and $kT$ is the effective blackbody temperature. Then one can solve the self-absorption frequency for weak absorption $\nu_{a}<{\rm min}(\nu_{m},\nu_{c})$ (intersection in the $1/3$ segment), intermediate absorption $\nu_{m}<\nu_{a}<\nu_{c}$ (intersection in the (p1)/2-(p-1)/2 segment) and strong absorption $\nu_{c}<\nu_{a}$ (electron pile-up around $\gamma_{a}$)
@@
where $\Sigma_{\rm p,tot}=\frac{N_{\rm p,tot}}{4\pi r^{2}}$ is the column density of the shocked proton and $N_{\rm p,tot}$ is the total number of shocked protons. The coefficient $0.92$ arises from the maximum value of the single-electron synchrotron spectrum function $\mathcal{F}(x)$ (Rybicki and Lightman, 1979), and the factor $\frac{\pi}{4}$ comes from the average pitch angle $\langle\sin\alpha\rangle$. The blackbody method adopted here is shown to be consistent with the optical depth method (Shen and Zhang, 2009) implemented in PyFRS (Lei et al., 2016) and PyBlastAfterglow (Nedora et al., 2024).
@@ 4.6 Synchrotron spectrum
Figure 8 shows an example of on-axis observed synchrotron spectra for narrow opening angle top-hat jets (thus mimicking the shape of the intrinsic spectrum) with afterglow parameters in the six different regimes.
@@ 4.7 Inverse Compton Cooling
[…]
@@
for slow cooling. The broken power-law/broken power-law division in Equation 119 and 120 give rise to additional spectral segments above $\gamma_{c}$ ($\nu_{c}$). As shown in Figure 8, the blue lines correspond to synchrotron-only cooling, the green lines to Thomson IC cooling with constant $\tilde{Y}$, and the orange lines to spectra including Klein–Nishina corrections, which introduce additional breaks for $\nu>\nu_{c}$.
@@ 4.8 Inverse Compton Spectrum
$I_{\nu}^{\rm IC}(\nu)$ $=$ $\delta^{\prime}\int_{1}^{\infty}\frac{dn_{e}}{d\gamma}d\gamma\int_{0}^{x_{0}}\sigma_{c}(x)I_{\nu_{0}}^{\rm syn}(x)dx$
$=$ $\int_{1}^{\infty}\frac{d\Sigma_{e}}{d\gamma}d\gamma\int_{0}^{\frac{\nu}{4\gamma^{2}x_{0}}}\sigma_{c}(\frac{\gamma h\nu_{0}}{m_{e}c^{2}})I_{\nu_{0}}^{\rm syn}(\nu_{0})\frac{\nu d\nu_{0}}{4\gamma^{2}\nu_{0}^{2}}\,$,
@@ 5 Geometry
$F_{\nu}(\nu,\tilde{t})=\frac{1+z}{4\pi D_{L}^{2}}\oiint_{t_{\rm obs}=\tilde{t}}4\pi r^{2}I_{\nu^{\prime}}(\nu^{\prime})\mathcal{D}^{3}d\Omega dr\,$, (132)
@@ 6 Examples
Medium type ISM Arbitrary ISM/Wind ISM/Wind Arbitrary
@@ 6.1 Strctured Jets
First, we present examples of structured jets observed from various viewing angles and compare them with the results from AfterglowPy and Jetsimpy (Wang et al., 2024). The left panel of Figure 10 shows the light curve for a top-hat jet with a half-opening angle of $5^{\circ}$. All three codes agree closely in this simple case; however, AfterglowPy exhibits a slight deviation at small viewing angles during the early coasting phase due to its approximate treatment of that regime. The middle panel illustrates a Gaussian jet with a $5^{\circ}$ core angle, while the right panel shows a power-law jet with index 2 and the same core angle. For structured jets, both AfterglowPy and Jetsimpy predict higher fluxes. This arises from their different treatments of the angular Lorentz‐factor profile: AfterglowPy allows $\Gamma(\theta)$ to fall below unity but truncates emission beyond a “wing” angle $\theta_{\rm wing}$, whereas Jetsimpy and VegasAfterglow model $\Gamma-1$ with a Gaussian profile and apply the “+1” correction.